OpenStudy (anonymous):

Write an exponential equation in the form y=ab^x which graph passes through (-3,24) (-2,12) I think its 3(24)^x

OpenStudy (freckles):

did you plug in the points into y=ab^x ? 

OpenStudy (freckles):

\[24=ab^{-3} \\ 12=ab^{-2}\]

OpenStudy (freckles):

this is a system of equations we have to solve

OpenStudy (freckles):

Like you could solve on of them for a in then use that in the other equation to solve for b

OpenStudy (anonymous):

I didn't understand any of that. I figured it out. thanks though.

OpenStudy (freckles):

\[25 =ab^{-3} \rightarrow 25b^{3}=ab^{-3}b^{3} \rightarrow 25b^3=ab^{0} \\ \rightarrow 25b^3=a(1) \rightarrow 25b^3=a\] see this first equation was solved for a

OpenStudy (freckles):

you are going to plug that into the other equation

OpenStudy (freckles):

and solve for b

OpenStudy (anonymous):

you i got it already. thanks!

OpenStudy (freckles):

oops 25 is suppose to be 24

OpenStudy (freckles):

ok